Please help me solve...
A women has $7000,part is invested for one year paying 7% interest and the remaining amount in a fund paying 9%. After a year total interest on the 2 investments is $540. How much is invested at each rate?

Answer by Mathtut(3670)

let amounts invested at 7% and 9% be x and y respectively
:
x+y=7000...........eq 1
.07x+.09y=540......eq 2
:
rewrite eq 1 to x=7000-y and plug this into eq 2
:
.07(7000-y)+.09y=540
:
490-.07y+.09y=540
:
.02y=50
:
y=2500 dollars invested at 9%
:
x=7000-y=7000-2500=4500 dollars invested at 7%